OpenStudy (anonymous):

Are the vectors s=(3, -4) and t=(11, 8) normal? Please help!

OpenStudy (anonymous):

@hartnn

OpenStudy (johnweldon1993):

Normal to a plane? or just normal in general?

OpenStudy (anonymous):

Just normal in general I guess

OpenStudy (johnweldon1993):

Well normal referring to a plane..means a vector perpendicular to the plane... So we need to see if these 2 vectors are perpendicular to each other... We can do that by calculating the dot product, and seeing if it = 0 If it is...then we have perpendicular vectors

OpenStudy (johnweldon1993):

Do you know how to calculate the dot product of 2 vectors?

OpenStudy (anonymous):

I am not sure. Do I just multiply x*x and y*y?

OpenStudy (anonymous):

Like 3*11 and -4*8?

OpenStudy (johnweldon1993):

Yeah and then add the results. \[\large <a,b>\cdot <c,d> = (a \times c) + (b \times d)\]

OpenStudy (anonymous):

So it equals to 1.

OpenStudy (johnweldon1993):

Correct.....and since that isn't 0 ...we find that these are not "normal" "perpendicular" vectors

OpenStudy (anonymous):

That's it? That was easy!

OpenStudy (johnweldon1993):

Exactly :) lol

OpenStudy (anonymous):

Lol thanks for your help! :)

OpenStudy (johnweldon1993):

Anytime :)